JoelTheBoss (joel_the_boss):

Given the function f(x) = x^3 + x^2 − 2x + 1, what is the resulting function when f(x) is shifted to the left 1 unit?

OpenStudy (anonymous):

when you shift a function b (number of) units to the left, you are essentially taking f(x+b).

JoelTheBoss (joel_the_boss):

I have no idea man.

OpenStudy (anonymous):

so for b units to the left you would be getting: g(x)=(x+b)^3+(x+b)^2-2(x+b)+1

OpenStudy (anonymous):

now, your b, that you are shifting to the left by, is 1.

JoelTheBoss (joel_the_boss):

Hmm.... How did you do that? lol Explain?

OpenStudy (anonymous):

sure, want to know why adding (x\(\normalsize\color{slate}{ \tt +b }\)) means a shift to the left, or just want to see what I did in this situation?

JoelTheBoss (joel_the_boss):

why when you add (x+b) it shifts to the left? Is it because its positive?

OpenStudy (anonymous):

was thinking about where to start. lets do this. if you shift a line you shift it up and left, right? 

OpenStudy (anonymous):

so see as you shifted the line up by "a" units, you have also shifted it to the left.... that is 1 reason. because by a line, when you add y=(x+a) inside the parenthesis, and add y=(x)+a outside the parenthesis, you are getting that same function, y=x+5. even though (x+a) is a shift to the left, and (x)+a is a shift up (by a units, in both).

OpenStudy (anonymous):

this is a one reason....

JoelTheBoss (joel_the_boss):

OHHHHH!! Ok, I see what you mean! They explained this in my lesson but didn't focus on it to much. Thanks man!

OpenStudy (anonymous):

same way, you will see that when you shift the line down, you are also shifting it to the right...